OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4

OpenStudy (anonymous):

ok, so expand the terms and collect the like terms together and solve! 3x+3 = -2x+2 -4

OpenStudy (anonymous):

rearrange and solve for x!

OpenStudy (anonymous):

-5?

OpenStudy (anonymous):

no, u ll have 3x + 2x = -3+2-4 which is 5x = -5 and x will be?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

Thank You (:

OpenStudy (anonymous):

:)